Diffusion Morning Report 06/28/11

Ted Berg tberg at stanford.edu
Tue Jun 28 07:46:23 PDT 2011


Good Morning all,
     Looking good this morning. Dome for epi is installed and cycle 
purging(we borrowed one from the new kit). TEOS had pressure 
issues-checking it out. Thermco 2 no TLC flow. Thermco 1 low growth. CDA 
to the fab was off this morning so some tools may have been down. No 
other issues at this time.

					
	Date: 	6/28/11 	THICKNESS-Change at 		Priority
Tube Name/Number 	Status 				
Tylan 1 	*up* 	** 	** 	** 	**
Tylan 2 	*up* 	** 	** 	** 	**
Tylan 3 	*up* 	** 	** 	** 	**
Tylan 4 	*up* 	** 	** 	** 	**
Tylan 5-BBr3 	*up* 	** 	** 	** 	**
Tylan 6-POCl3 	*up* 	** 	** 	** 	**
Tylan 7-TEOS2 	*yellow* 	*pressure issues* 	*10.14/30* 	** 	*2*
Tylan 8-SiGe 	*up* 	** 	*1.94/30* 	** 	**
Tylan 9-Poly 	*up* 	** 	*22.3/30* 	** 	**
Tylan 10-Nitride 	*up* 	** 	*16.51/30* 	** 	**
Tylan 11-LTO 	*up* 	** 	*10.4/30* 	** 	**
Tylan 12-FGA 	*up* 	** 	** 	** 	**
Thermco 1 	*yellow* 	*low growth rate* 	** 	** 	**
Thermco 2 	*yellow* 	*no tlc flow* 	** 	** 	**
Thermco 3 	*future* 	** 	** 	** 	**
Thermco4 	*future* 	** 	** 	** 	**
Thermco Poly1 	*up* 	*
* 	** 	** 	**
	*up* 	** 	** 	** 	**
Thermco Nitride1 	*yellow* 	*pressure issues* 	** 	** 	**
Thermco bpsg 	*up* 	** 	** 	** 	**
AMAT Epi 	*down* 	*dome installed-cycle purging* 	** 	** 	*1*
TEL Spa 	*up* 	** 	** 	** 	**
Burn Box 	*up* 	** 	** 	** 	**
Ebeam 	*down* 	*cryo compresssor dead* 	** 	** 	**
ZYGO 	*up* 	*
* 	** 	** 	**
FGA2 	*up* 	** 	** 	** 	**
SiH4 for furnaces 	** 	*B on line pressure 938 psi* 	*changeover working 
well* 	** 	**




-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20110628/e54fd4e0/attachment.html>


More information about the diffusion mailing list